OpenStudy (anonymous):

Derivative of (ln^2)x

OpenStudy (anonymous):

\[\frac{ 2 \ln(x) }{ x }\]

OpenStudy (mathstudent55):

What kind of notation is ln^2?

OpenStudy (anonymous):

\[\ln ^2 (x) = \ln(x) \times \ln(x)\]

OpenStudy (anonymous):

\[f(x) = \ln^2(x)\] \[f(x)=\ln (x) \times \ln(x)\] \[f'(x) = \ln(x) \times \frac{ 1 }{ x } + \frac{ 1 }{ x }\times \ln(x)\] \[f'(x) = \frac{ \ln(x) }{ x } + \frac{ \ln(x) }{ x }\]

OpenStudy (mathstudent55):

Thanks. \( \dfrac{d}{dt} (\ln x)^2 \) \(= 2 \ln x \dfrac{d}{dt} (\ln x) \) \(= 2 \ln x \cdot \dfrac{1}{x} \) \(= \dfrac{ 2 \ln x }{x} \)